How many grams of CaO are required to neutralize 852 gm of P410? Draw structure of P410 molecule.

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
6CaO + P4O10 → 2Ca3 (PO4)2
Moles of P4­O10 = 852/284 = 3
Moles of CaO = 3 * 6 = 18
Wt. of CaO = 18 * 56 = 1008 g
For structure of P4O10 : See question 20 of this Section.